OpenStudy (anonymous):

Which of the following is equivalent to log 4 32 ?

OpenStudy (across):

I believe the "Tweet this Question" button is more equivalent to \(\log_432\) than the "Share this Question" button.

OpenStudy (callisto):

\[\log_432\]= log32/log 4 = log (2^5) / log (2^2) =(5log2) / (2log2) = 5/2